# On the best approximation by finite Gaussian mixtures

### v1

@@ 5 Applications and extensions
In this section, we discuss a few applications and extensions of the paper.
@@ 5.1 Convergence rates of nonparametric maximum likelihood estimator
$\hat{P}\in\mathop{\arg\max}_{Q\in{\mathcal{Q}}}\frac{1}{n}\sum_{i=1}^{n}\log f_{Q}(X_{i}),$ (28)
where the optimization is over a class ${\mathcal{Q}}$ of priors. The finite-order approximation of a family of distributions naturally yields upper bounds of the metric entropy of that family, which further determines the convergence rate of the NPMLE [WS95, vdVW96, GvdV01, JZ09]. Following the analysis in [Zha09], the Hellinger distance $H(f_{\hat{P}},f_{P^{\star}})$ can be upper bounded in terms of $m^{\star}(\epsilon,{\mathcal{Q}},L_{\infty}[-M,M])$, where $M$ is a parameter depending on the tail of ${P^{\star}}$. Hence, our analysis can recover the convergence rates derived in [Zha09].
Moreover, suppose $P^{\star}$ belongs to a given family of distributions ${\mathcal{P}}$, Theorem 8 below in fact establishes a faster rate for the constrained NPMLE with ${\mathcal{Q}}={\mathcal{P}}$, which is characterized by $m^{\star}(H,{\mathcal{P}},n^{-1/2})\frac{\log n}{n}$ with respect to the Hellinger distance.444 We note that existing minimax lower bounds in [PW21, Theorem 20] agree with $m^{\star}(H,{\mathcal{P}},n^{-1/2})/n$. However, bridging this gap remains an open problem. The improvement arises from our tight approximation rates. In contrast to [Zha09, Theorem 1] which yield the same convergence rates for the families ${\mathcal{P}}^{\textsf{Bdd}}_{M}$ and ${\mathcal{P}}_{\alpha}(\beta)$, our result gives a strictly faster rate for the former. In addition, another technical contribution is a a simplified proof: Compared with that of [Zha09, Theorem 1], we simplify the argument by first proving a $L_{\infty}$-entropy estimate then directly bounding the square root of the mixture density. The proofs are deferred to Appendix D.
@@
    Let $P^{\star}\in{\mathcal{P}}^{\textsf{Bdd}}_{M}$, where $M\leq n^{c}$ for some constant $c>0$. Let $\hat{P}$ be a solution of (28) with ${\mathcal{Q}}={\mathcal{P}}^{\textsf{Bdd}}_{M}$. There exist constants $s^{\star},c^{\prime}>0$ such that for any $s\geq s^{\star}$,
@@
    Let $P^{\star}\in{\mathcal{P}}_{\alpha}(\beta)$, where $\alpha>0$ is fixed and $\beta\leq n^{c}$ for some constant $c>0$. Let $\hat{P}$ be a solution of (28) with ${\mathcal{Q}}={\mathcal{P}}_{\alpha}(\beta)$. There exist $s^{\star},c^{\prime}>0$ which depend on $\alpha$ such that for any $s\geq s^{\star}$, (29) holds with
@@ 5.2 Mixing distributions under moment constraints
So far, our results deal with mixing distributions that exhibit exponential decay in their tails. In fact, our analysis framework also allows for extension to distributions with heavier tails. Let us consider mixing distributions under moment conditions, which were well-studied in mixture models [GvdV07b, Zha09, SG20, SW23] Let $\psi(x)=x^{\alpha}$ for $\alpha>0$. Then the corresponding $\psi$-Orlicz norm is simply the usual $L_{\alpha}$-norm $\|X\|_{\psi}=\|X\|_{\alpha}\triangleq\mathbb{E}[|X|^{\alpha}]^{\frac{1}{\alpha}}$. Define
@@ 5.3 Generalization to other mixture models
So far we have focused on approximating general one-dimensional Gaussian location mixture by finite mixtures. Our techniques of finite-order approximation can also adapt to other models, including the Gaussian location-scale mixtures and multi-dimensional mixtures. In these contexts, most of the existing approximation results [SG20, Jia20, SGS21, SW23] in the empirical Bayes literature rely on Taylor expansion and local moment matching.
Location-scale Gaussian mixtures
@@
[…]
@@
Previous works applied moment-based approach to construct approximators. For instance, when $H$ is supported in $[-1,1]^{2}$, by applying Taylor expansion and Carathéodory’s theorem, [GvdV01, Lemma 3.4] constructs an $H_{m}$ with $m\leq O(\log^{2}\frac{1}{\epsilon})$ such that $\|{f_{H}-f_{H_{m}}}\|_{\infty}\leq\epsilon$. In addition, if $X$ and $S$ are independent, both $X_{m}$ and $S_{m}$ are supported on $O(\log\frac{1}{\epsilon})$ points.
@@
It is unclear how to lower bound this uniformly over $(X_{m},H_{m})$ that is $m$-atomic. However, for the special case of scale mixtures with $X\equiv 0$ and $S$ bounded away from zero and infinity, we can further lower bound (32) as follows. First, by choosing $\omega\in\{\sqrt{k}:k\in\{0,\ldots,2m\}\}$, using the same argument as in Proposition 2, (32) can be related to the minimum eigenvalue $\lambda_{\min}$ of the Hankel matrix $\left(\mathbb{E}\exp\left(-\frac{(j+k)S^{2}}{2}\right)\right)_{j,k=0}^{m}$. Applying the method based on orthogonal expansion in Proposition 3 (cf. Appendix B.2), one can show that $\lambda_{\min}\geq e^{-\Omega(m)}$, leading to the lower bound that at least $\Omega(\log\frac{1}{\epsilon})$ components are needed to achieve an approximation error $\epsilon$ in the scale mixture model. Establishing lower bounds for general location-scale mixtures remains an outstanding challenging.
Multiple dimensions
Consider the Gaussian location mixture on $\mathbb{R}^{d}$:
@@
The lower bound is derived following the roadmap in Section 4.1. For the multidimensional regime, the trigonometric moment matrices are replaced by multilevel Toeplitz matrices [TZ98, Pes19], and the wrapped density approach is then applied to yield the following result.
@@
${\mathcal{E}}^{\star}(m,\mathrm{Unif}([-M/\sqrt{d},M/\sqrt{d}]^{d}),\mathsf{TV})\geq\frac{1}{2^{d+1}m}\exp\left(-\frac{d^{2}\pi^{2}m^{\frac{2}{d}}}{2M^{2}}\right),$
The approximation rates for the distribution classes under tail conditions can be similarly extended. The additional results and proofs are presented in Appendix C.3.

### v2

@@ 5 Applications and extensions
In this section, we discuss the applications of our results in information theory and nonparametric statistics. Additionally, we show how our findings can be extended to a wider range of distribution families and mixture models.
5.1 Constellation design for Gaussian channels
The finite mixture approximation problem is closely related to the input constellation design for Gaussian channels. Recall that the Gaussian distribution is capacity-achieving for Gaussian channels under average power constraint. Under an additional input cardinality constraint of at most $m$, the capacity gap is characterized by ${\mathcal{E}}^{\star}(m,N(0,\sigma^{2}),\mathsf{KL})$ (see (3)). Consequently, our proposed approximator serves as an $m$-point constellation that approaches the limit at the optimal rate.
We implement our finite mixture approximation in Algorithm 1, utilizing the moment-matching approach for truncated distributions as described in Section 3.2. Theorem 4 implies that setting $t\asymp\sigma\sqrt{m\log(1+\sigma^{-1})}$ and $K\asymp\sigma^{2}\log(1+\sigma^{-1})$ yields a rate-optimal approximator. For the Gauss quadrature, we compute the moments of conditional distributions and apply the moment-based Golub-Welsch algorithm [GW69], as detailed in Algorithm 2 in Appendix A.2.
Algorithm 1 TruncQuad constellation for Gaussian channels
  Input: Number of atoms $m$, signal-noise ratio $\sigma^{2}$, $c_{1},c_{2},c_{3}>0$
  Output: $m$-atomic distribution $P_{m}\in{\mathcal{P}}_{m}$
  Define $P=N(0,\sigma^{2})$, $t=c_{1}\sigma\sqrt{m\log(c_{2}+\sigma^{-1})}$, $K=\min\left(\lceil c_{3}t^{2}/m\rceil,m\right)$, and $m_{0}=\lfloor m/K\rfloor$
  Partition the intervals: $I_{j}=[-t+(j-1)\frac{2t}{K},-t+j\frac{2t}{K}],j\in[K]$
  Compute the conditional moment $m_{j}(P_{(k)})=\int_{I_{k}}x^{j}\mathrm{d}P/P(I_{k})$ for $j\in[2m_{0}-1]$ and $k\in[K]$
  For each $k\in[K]$, compute $P_{k}$ by Algorithm 2 with input ${\mathbf{m}}_{k}=\left(m_{j}(P_{(k)})\right)_{j=1}^{2m_{0}-1}$
  Compute $P_{m}=\sum_{k=1}^{K}P(I_{k})P_{k}$
We conduct numerical experiments to evaluate the capacity gap of the truncated quadrature method (TruncQuad) described in Algorithm 1 with parameters $(c_{1},c_{2},c_{3})=(2,5,0.2)$. Fixing $\sigma^{2}\in\{1,10,50\}$, we measure the capacity gap as the number of atoms $m$ increases. We compare its performance against the following constellation schemes considered in [WV10b]:
  1. (1)
    Equilattice: the constellation uniformly distributed on the grid $E_{m}$, defined as
    $E_{m}=\begin{cases}\left\{2i\Delta_{m}:i=\frac{1-m}{2},\ldots,0,\ldots,\frac{m-1}{2}\right\}&m\text{ odd }\\ \left\{(2i+1)\Delta_{m}:i=-\frac{m}{2},\ldots,0,\ldots,\frac{m-2}{2}\right\}&m\text{ even, }\end{cases}\quad\Delta_{m}=\sqrt{\frac{3}{m^{2}-1}}.$
  2. (2)
    Quadrature: the Gauss quadrature rule based on the first $12m-$ 1 moments of $P$.
  3. (3)
    Quantized: the real line is partitioned into $m$ subintervals using quantile sequences $\alpha_{j}=\Phi^{-1}\left(\frac{j-1}{m}\right)$, $j=0,\ldots,m$. The $m$-point uniform distribution is then supported on the conditional expectations $x_{j}=\mathbb{E}_{P}[X\mid X\in[\alpha_{j-1},\alpha_{j}]]$ for $j\in[m]$, with the constellation scaled to achieve variance $\sigma^{2}$.
  4. (4)
    Random walk: sample $B_{m}\sim\operatorname{Binomial}(m-1,1/2)$ and define $X_{m}=\frac{2\sigma}{\sqrt{m-1}}\left(B_{m}-\frac{m-1}{2}\right)$ as the input. By the central limit theorem, $\hat{X}_{m}$ asymptotically converges to $N(0,\sigma^{2})$.
Figure 1 presents the results, showing that both TruncQuad and Gauss-Hermite quadrature converges exponentially, whereas the quantized Gaussian and random walk schemes converge more slowly at a polynomial rate. Meanwhile, the equilattice input achieves only a constant error, as it converges weakly to $\mathrm{Unif}([-\sqrt{3},\sqrt{3}])$ as $m\to\infty$. Moreover, TruncQuad yields smaller errors than Gauss-Hermite quadrature and significantly so for high SNR (large $\sigma$), which aligns with our theoretical analysis. Notably, TruncQuad converges at a rate of $\exp\left(-\Omega\left(m\log\left(1+\frac{1}{\sigma}\right)\right)\right)$ (see (17)), outperforming the $\sigma^{2}\exp\left(-\Omega\left(m\log\left(1+\frac{1}{\sigma^{2}}\right)\right)\right)$ rate of Gauss-Hermite quadrature [WV10b, Theorem 8] when $\sigma$ is large.
Figure 1: Capacity gap with $\sigma^{2}=1,10,50$ as $m$ increases.
@@ 5.2 Convergence rates of nonparametric maximum likelihood estimator
${\hat{P}({\mathcal{Q}})\in\mathop{\arg\max}_{Q\in{\mathcal{Q}}}\frac{1}{n}\sum_{i=1}^{n}\log f_{Q}(X_{i}),}$ (31)
where the optimization is over a class ${\mathcal{Q}}$ of priors. The finite-order approximation of a family of distributions naturally yields upper bounds of the metric entropy of that family, which further determines the convergence rate of the NPMLE [WS95, vdVW96, GW00, GvdV01, JZ09]. Particularly, [Zha09] studies the unconstrained NPMLE where ${\mathcal{Q}}$ is the class of all distributions on $\mathbb{R}$. Following the analysis in [Zha09], the Hellinger distance $H(f_{\hat{P}},f_{P^{\star}})$ can be upper bounded in terms of $m^{\star}(\epsilon,{\mathcal{Q}},L_{\infty}[-M,M])$. In this regard, our analysis recovers the convergence rates established in [Zha09] by deriving an upper bound for $m^{\star}(\epsilon,\mathcal{Q},L_{\infty}[-M,M])$ as an application of Theorem 3. This result is detailed in Lemma 12.
Moreover, suppose $P^{\star}$ belongs to a given family of distributions ${\mathcal{P}}$, Theorem 8 below in fact establishes a faster rate for the constrained NPMLE with ${\mathcal{Q}}={\mathcal{P}}$, which is characterized by $m^{\star}(n^{-1/2},{\mathcal{P}},H)\frac{\log n}{n}$ with respect to the Hellinger distance333 We note that existing minimax lower bounds in [PW21, Theorem 20] agree with $m^{\star}(n^{-1/2},{\mathcal{P}},H)/n$. However, bridging this gap remains an open problem.. In contrast to [Zha09, Theorem 1] which yield the same convergence rates of the unconstrained NPMLE for the families ${\mathcal{P}}^{\textsf{Bdd}}_{M}$ and ${\mathcal{P}}_{\alpha}(\beta)$, our result gives a strictly faster rate for the former. The improvement stems from the additional structural information $P^{\star}\in\mathcal{P}$, which enables the application of our tight approximation rates. In addition, another technical contribution is a simplified proof: Compared with that of [Zha09, Theorem 1], we simplify the argument by first proving a $L_{\infty}$-entropy estimate then directly bounding the square root of the mixture density. The proofs are deferred to Appendix D.
@@
    Let $P^{\star}\in{\mathcal{P}}^{\textsf{Bdd}}_{M}$, where $M\leq n^{c}$ for some constant $c>0$. Denote $\hat{P}=\hat{P}({\mathcal{P}}^{\textsf{Bdd}}_{M})$. There exist constants $s^{\star},c^{\prime}>0$ such that for any $s\geq s^{\star}$,
@@
    Let $P^{\star}\in{\mathcal{P}}_{\alpha}(\beta)$, where $\alpha>0$ is fixed and $\beta\leq n^{c}$ for some constant $c>0$. Denote $\hat{P}=\hat{P}({\mathcal{P}}_{\alpha}(\beta))$. There exist $s^{\star},c^{\prime}>0$ which depend on $\alpha$ such that for any $s\geq s^{\star}$, (32) holds with
@@
Remark 5.
Theorem 8 can be extended to other divergences using the divergence comparison inequalities [JPW23]. For example, for $d=\mathsf{KL}$ and $\chi^{2}$ and any $P,Q\in{\mathcal{P}}^{\textsf{Bdd}}_{M}$, the inequalities take the form $d(f_{P},f_{Q})\leq C_{M}H^{2}(f_{P},f_{Q})$ with some $C_{M}$ depending on $M$. By (32), we have
$\mathbb{P}[d(f_{\hat{P}},f_{P^{\star}})\geq s^{2}C_{M}\epsilon_{n}^{2}]\leq\mathbb{P}[H(f_{\hat{P}},f_{P^{\star}})\geq s\epsilon_{n}]\leq n^{-c^{\prime}s^{2}},$
which implies an $O(C_{M}\epsilon_{n}^{2})$ rate in $d(f_{\hat{P}},f_{P^{\star}})$ for the constrained NPMLE on ${\mathcal{P}}^{\textsf{Bdd}}_{M}$. The tight dependency on the parameters $M$ remains open.
@@ 5.3 Mixing distributions under moment constraints
So far, our results deal with mixing distributions that exhibit exponential decay in their tails. In fact, our analysis framework also allows for extension to distributions with heavier tails. Let us consider mixing distributions under moment conditions, which were well-studied in mixture models [GvdV07, Zha09, SG20, SW24] Let $\psi(x)=x^{\alpha}$ for $\alpha>0$. Then the corresponding $\psi$-Orlicz norm is simply the usual $L_{\alpha}$-norm $\|X\|_{\psi}=\|X\|_{\alpha}\triangleq\mathbb{E}[|X|^{\alpha}]^{\frac{1}{\alpha}}$. Define
@@ 5.4 Generalization to other mixture models
So far we have focused on approximating general one-dimensional Gaussian location mixture by finite mixtures. Our techniques of finite-order approximation can also adapt to other models, including the Gaussian location-scale mixtures and multi-dimensional mixtures. In these contexts, most of the existing approximation results [SG20, Jia20, SGS24, SW24] in the empirical Bayes literature rely on Taylor expansion and local moment matching.
Location-scale Gaussian mixtures.
@@
$f_{H}(x)=\int\frac{1}{\sigma}\phi\left(\frac{x-\theta}{\sigma}\right)\mathrm{d}H(\theta,\sigma),$
@@
Previous works applied moment-based approach to construct approximators. Suppose that $H$ is supported on $[-a,a]\times[\underline{\sigma},\bar{\sigma}]$ for some positive constants $a$ and $\underline{\sigma}\leq\bar{\sigma}$. By applying Taylor expansion and Carathéodory’s theorem, [GvdV01, Lemma 3.4] constructs an $H_{m}$ with $m\leq O(\log^{2}\frac{1}{\epsilon})$ such that $\|{f_{H}-f_{H_{m}}}\|_{\infty}\leq\epsilon$. In addition, if $X$ and $S$ are independent, both $X_{m}$ and $S_{m}$ are supported on $O(\log\frac{1}{\epsilon})$ points.
Under the above compact support condition, a location-scale Gaussian mixture can be reduced to a location mixture. Specifically, define $X^{\prime}=X+\sqrt{S^{2}-\underline{\sigma}^{2}}Z\sim P^{\prime}$ and $Z_{\underline{\sigma}}\sim N(0,\underline{\sigma}^{2})$. Then
$X^{\prime}+Z_{\underline{\sigma}}\sim f_{H},$
and $X^{\prime}$ is $\sigma^{2}$-subgaussian for some $\sigma$ depending on $a$ and $\bar{\sigma}$. Applying (8) and Remark 1, there exists an $m$-atomic $P_{m}$ on $\mathbb{R}$ with $m\leq O(\log\frac{1}{\epsilon})$ such that $\mathsf{TV}(f_{P_{m}},f_{H})\leq\epsilon$. Furthermore, since $\|f_{H}^{\prime}\|_{\infty}\leq O(1)$, $\mathsf{TV}(f_{P_{m}},f_{H})\lesssim\epsilon$ implies $\|{f_{P_{m}}-f_{H}}\|_{\infty}\lesssim\sqrt{\epsilon}$. Consequently, the $O(\log\frac{1}{\epsilon})$ upper bound also holds for $d=L_{\infty}$, thereby improving the previous results in [GvdV01].
@@
It is unclear how to lower bound this uniformly over $(X_{m},H_{m})$ that is $m$-atomic.
A key challenge arises from the distinct effects of the location and scale parameters on the mixture distribution. Following our framework, the corresponding low-rank matrix over a grid $\omega=\{\omega_{j}\}_{j=0}^{m}\subseteq\mathbb{R}$ can be constructed as
$\mathbf{V}_{m}\triangleq\mathbb{E}\left[\mathbf{v}_{m}(X_{m},S_{m})\mathbf{v}_{m}^{\star}(X_{m},S_{m})\right]=\left(\mathbb{E}\left[e^{-i(\omega_{j}-\omega_{k})X_{m}-\frac{1}{2}(\omega_{j}^{2}+\omega_{k}^{2})S_{m}^{2}}\right]\right)_{j,k=0}^{m},$
where $\mathbf{v}_{m}(x,s)\triangleq\left(1,e^{-i\omega_{1}x-\frac{1}{2}\omega_{1}^{2}s^{2}},\dots,e^{-i\omega_{m}x-\frac{1}{2}\omega_{m}^{2}s^{2}}\right)$. However, the entries in $\mathbf{V}_{m}$ differ from the terms in the variational lower bound (35).
Nevertheless, for the special case of scale mixtures with the additional constraint $X_{m}=X\equiv 0$, we can further lower bound (35) as follows. First, by choosing $\omega\in\{\sqrt{k}:k\in\{0,\ldots,2m\}\}$, using the same argument as in Proposition 2, (35) can be related to the minimum eigenvalue $\lambda_{\min}$ of the Hankel matrix $\left(\mathbb{E}\exp\left(-\frac{(j+k)S^{2}}{2}\right)\right)_{j,k=0}^{m}$. Applying the method based on orthogonal expansion in Proposition 3 (cf. Appendix B.2), one can show that $\lambda_{\min}\geq e^{-\Omega(m)}$, leading to the matching lower bound that at least $\Omega(\log\frac{1}{\epsilon})$ components are needed to achieve an approximation error $\epsilon$ under $\mathsf{TV}$ in the scale mixture model. Establishing lower bounds for general location-scale mixtures remains an outstanding challenging.
Multiple dimensions.
Fix $d\in\mathbb{N}$. Consider the Gaussian location mixture on $\mathbb{R}^{d}$:
@@
The lower bound is derived following the roadmap in Section 4.1. For the multidimensional regime, the trigonometric moment matrices are replaced by multilevel Toeplitz matrices [TZ98, Pes19], and both the wrapped density approach and the orthogonal expansion approach are then applied to yield the following result.
@@
${\mathcal{E}}^{\star}(m,\mathrm{Unif}([-M/\sqrt{d},M/\sqrt{d}]^{d}),\mathsf{TV})\geq\frac{1}{2^{d+1}m}\exp\left(-\frac{d^{2}\pi^{2}m^{\frac{2}{d}}}{2M^{2}}\right).$ (37)
Furthermore, there exists a universal constant $C_{1}$ and $C^{\prime}$ such that when $m^{\frac{1}{d}}\geq C^{\prime}M^{2}/d$,
${\mathcal{E}}^{\star}(m,{\mathcal{P}}^{\textsf{Bdd}}_{M,d},\mathsf{TV})\geq\frac{1}{2^{d+1}m}\exp\left(-C_{1}dm^{\frac{1}{d}}\log\left(\frac{dm^{\frac{1}{d}}}{M^{2}}\right)\right).$ (38)
Consequently, Propositions 5 and 6 give the tight approximation rate of $m$ and $M$ for the multi-dimensional distributions with compact support. The approximation rates for the distribution classes under tail conditions can be similarly extended. The additional results and proofs are presented in Appendix C.3.
General mixtures.
Our analytical framework extends beyond Gaussian mixtures. For the lower bounds, the convolution $P*\nu$ is the distribution of $X+Z$ with $X\sim P$ and $Z\sim\nu$ being independent. Analogous to Proposition 2, we obtain
$\inf_{P_{m}\in{\mathcal{P}}_{m}}\mathsf{TV}(P_{m}*\nu,P*\nu)\geq\sup_{\delta>0}\frac{\lambda_{\min}({\mathbf{T}}_{m}(\delta X))}{2(m+1)\max_{k\in\mathbb{Z},|k|\leq m}|t_{k}(\delta Z)|}.$
The lower bound for approximating $P*\nu$ thus follows from bounding $\lambda_{\min}(\mathbf{T}_{m}(\delta X))$, as discussed in Section 4.1.2, together with the formula for $t_{k}(\delta Z)$.
The upper bounds can also be extended if the mixture can be identified by sufficiently many moments of its mixing distribution, akin to Lemma 1. For instance, consider approximating the Poisson mixture
$\pi_{P}(y)\triangleq\int\operatorname{poi}(y,\lambda)\mathrm{d}P(\lambda)$
by some finite mixture $\pi_{P_{m}}$, where $\operatorname{poi}(y,\lambda)=\frac{e^{-\lambda}\theta^{y}}{y!}$ and $P_{m}\in{\mathcal{P}}_{m}$. As an Poisson analogue of Lemma 1, [WY20b, Theorem 3.3.4] shows that for any $P,Q$ supported on $[0,M]$, if $m_{j}(P)=m_{j}(Q)$ for all $j\in[L]$, then
$\mathsf{TV}\left(\pi_{P},\pi_{Q}\right)\leq\left(\frac{eM}{2L}\right)^{L}.$
For $P\in{\mathcal{P}}^{\textsf{Bdd}}_{M}$ with $M\lesssim\log\frac{1}{\epsilon}$, there exists $P_{m}\in{\mathcal{P}}_{m}$ with $m\leq O\left(\frac{\log\frac{1}{\epsilon}}{\log\left(1+\frac{1}{M}{\log\frac{1}{\epsilon}}\right)}\right)$ such that $\mathsf{TV}\left(\pi_{P_{m}},\pi_{P}\right)\leq\epsilon$. For larger $M$, [SW24, Lemma 6] suggests that the result may be further improved by matching the moments of conditional distributions on quadratically partitioned subintervals, and the optimal rate remains open.